# Assessing the star formation history of all-sky and part-sky 100 pc white dwarf samples

### v1

@@ Abstract
Thanks to Gaia and large-scale spectroscopic follow-up surveys (4MOST, DESI, WEAVE, SDSS-V), it is now possible to build representative and minimally biased samples of the local white dwarf population. Here we analyse several volume-limited 100 pc samples of white dwarfs, constructed from different surveys and studies, to evaluate their completeness and residual biases. We model the underlying star formation history and Galactic disc age via comparison with simulated populations of white dwarfs to quantitatively characterise completeness. We assess whether the benefit of Gaia XP spectra in datasets outweighs the reduction in sample size, and to what extent targeted, part-sky, and magnitude limited surveys can be used in comparison to all-sky volume limited surveys. Additionally, we simulate the 4MOST 100PC sub-survey and discuss its use to better understand the local star formation history.
@@
pubyear: 2025pagerange: Assessing the star formation history of all-sky and part-sky 100 pc white dwarf samplesDirect Age Calculations
@@ 1 Introduction
White dwarfs have long been recognised for their potential as chronometers, since they cool over time and have no source of nuclear fuel to maintain their temperatures (Mestel, 1952). With a close relationship between effective temperature, luminosity (or mass), and their total age (main-sequence lifetime plus white dwarf cooling time), they provide a way of rewinding time in a stellar population to look at the underlying star formation rates (Winget et al., 1987; Oswalt et al., 1996; Rowell, 2013; Tremblay et al., 2014; Isern, 2019; Fantin et al., 2019; Torres et al., 2021; Cukanovaite et al., 2023; Roberts et al., 2025). Realising this potential, however, requires detailed models of white dwarf cooling times, main sequence lifetimes, the initial mass function, the initial-final mass relation, and a careful treatment of observational biases. These variables contribute uncertainty and cloud our knowledge of the star formation history (SFH) that produced the white dwarfs. A common benchmark is to compare such reconstructions with the formation histories inferred from main-sequence stars in the same populations (Mor et al., 2019; Alzate et al., 2021; Dal Tio et al., 2021; Gallart et al., 2024; Fernández-Alvar et al., 2025), with the broader goal of informing models of Galactic formation and evolution.
@@
Several methods have been used to derive the SFH of a white dwarf population based on different observational properties: estimating individual total ages for each star or using stellar population synthesis models to fit the luminosity function, absolute magnitude distribution, or Hertzsprung–Russell diagram. In a recent study, Roberts et al. (2025) compared these approaches using the 40 pc Gaia-identified, volume-complete sample of 1073 spectroscopically confirmed white dwarfs (O’Brien et al., 2024). They found that, within systematic uncertainties and finite sample size, the different methods yielded broadly consistent results. Among them, the method based on the absolute Gaia G-band magnitude distribution was identified as the most promising, since it requires only a parallax and a G-band magnitude and can therefore be applied to larger and more distant samples with less complete follow-up data. Using these methods, they showed that a constant SFH provided a good overall fit to the sample, although variations by up to a factor of about two — either in the recent past (e.g. Mor et al., 2019) or at earlier times—could not be ruled out given the uncertainties in population modelling.
@@
With each successive Gaia data release (Gaia Collaboration et al., 2016, 2018, 2023), the size and precision of stellar samples available for Galactic studies has grown dramatically. The advent of machine learning classification, improved stellar models and quality filtering, and increasingly realistic Galactic simulations have further enhanced our ability to extract astrophysical information from these data. In parallel, large efforts have been devoted to the spectroscopic characterisation of Gaia catalogues: from the low-resolution XP spectrophotometry provided on board Gaia, to ground-based multi-object spectroscopic (MOS) surveys such as DESI (DESI Collaboration et al., 2025), 4MOST (de Jong et al., 2019), WEAVE (Jin et al., 2024), and SDSS-V (Kollmeier et al., 2025), which together provide the medium-resolution optical ($\approx 3600-10\,000$ Å) spectroscopic context essential for Galactic archaeology and beyond. This enables more precise determinations of stellar parameters, chemical abundances, radial velocities, magnetic field strengths, and binary properties that are not possible with Gaia alone.
@@
There now exist a range of 100 pc Gaia-defined samples of confirmed and candidate white dwarfs (Gentile Fusillo et al., 2021; Gaia Collaboration et al., 2021; Jiménez-Esteban et al., 2023; García-Zamora et al., 2023; Vincent et al., 2024), constructed with varying quality cuts, white dwarf probability thresholds, and each of them can be combined with auxiliary data, leading to differences in sky coverage. In particular, the footprint of Gaia-selected white dwarfs observed in MOS surveys currently covers only about 20% of the sky (DESI DR1; DESI Collaboration et al. 2025), with plans for each MOS mission to expand to roughly 50% of the sky by the end of operations. While these catalogues contain an order of magnitude more white dwarfs than the 40 pc sample, reducing Poisson errors on the SFH, and allowing for a wider probe of the Solar neighbourhood including the effect of the finite vertical scale height of the Galactic disc, they also lack certain things the 40 pc sample has such as either full volume completeness or spectroscopic follow up. Balancing these factors and assessing their impact is crucial for selecting the best white dwarf samples for population studies.
@@
In this work, based on the 100 pc Gaia stellar sample, we compare different white dwarf sample selections, either based on quality cuts or availability of auxiliary data such as XP spectrophotometry or MOS medium-resolution spectroscopy. We then rely on the stellar population code of Roberts et al. (2025) to quantify these differences in terms of changes in inferred SFH and age of the Galactic disc.
@@
Identifying white dwarfs in the Gaia DR3 catalogue is challenging, even within the relatively nearby 100 pc volume. As shown for example by Gentile Fusillo et al. (2021), contamination from non–white dwarf objects remains non-negligible, requiring probabilistic classification to separate white dwarfs from main-sequence stars, unresolved binaries and spurious sources. Nevertheless, the completeness of the Gaia 100 pc white dwarf sample is expected to be high, on the order of 93 per cent from a comparison with the SDSS footprint (Gentile Fusillo et al., 2021). This is in line with an independent estimate by O’Brien et al. (2023, 2024), which suggests that the Gaia white dwarf selection of Gentile Fusillo et al. (2021) achieves $\approx$97 per cent completeness in the 40 pc sample from a comparison with previously known white dwarfs not identified in Gaia, many of them members of unresolved binaries.
The sensitivity of the 100 pc selection is also demonstrated by different authors reaching different selections. The Gaia Collaboration’s Catalogue of Nearby Stars (Gaia Collaboration et al., 2021) (GCNS) finds 14 699 high-probability white dwarfs ($P_{\rm WD}\geq 0.75$) within 100 pc compared to the 15 001 found in Gentile Fusillo et al. (2021) for the same threshold in $P_{\rm WD}$. Only 14 687 of these sources are in common between both catalogues. Also, the independent analysis of Jiménez-Esteban et al. (2023) arrives at a different total of 12 718 white dwarfs within 100 pc. These reflect the different HR diagram selections, definition of white dwarf probability thresholds, and adopted quality cuts, underscoring the difficulty of assembling a complete volume census from Gaia alone.
Restricting the population to those white dwarfs with additional spectroscopic information reduces the available sample substantially. Limiting to objects with Gaia XP spectrophotometry, as explored by Vincent et al. (2024), results in 13 235 white dwarfs within 100 pc. Limiting to objects with existing ground-based spectroscopy within the SDSS footprint (Kilic et al., 2025) or DESI DR1 (DESI Collaboration et al., 2025) results in samples of 2 750 and 2 978 white dwarfs, respectively. While these cuts enhance the reliability of atmospheric parameter estimates and classifications, they inevitably reduce the sample size and may introduce biases. Most of this loss arises from the limited (primarily northern hemisphere) footprints of these MOS surveys, but within their surveyed conical 100 pc volumes the resulting samples remain reasonably complete and representative (Kilic et al., 2020; Manser et al., 2024).
@@
Table 1: Sample sizes and quality cuts made for the samples considered in this work. The GCNS catalogue (Gaia Collaboration et al., 2021) also includes main-sequence stars, hence the large number of objects removed after the $P_{\rm WD}$ white dwarf selection. The Jiménez-Esteban sample does not have masses and effective temperatures for all stars in the sample; where they are missing, the values from GF21 are used instead. The part-sky DESI and SDSS footprints are shown in Fig. 3. ‘tailored atm’ implies that photometric or spectroscopic fits were performed with model atmospheres using an appropriate chemical composition for a given known spectral type. In all cases, photometric atmospheric parameters were corrected for the low-mass problem at $T_{\rm eff}<6000$ K (O’Brien et al., 2024) before the mass cuts for binary removal were performed. Note that in the absolute $G$ magnitude distribution method, we do not use the derived stellar parameters.
@@
Stars in
catalogue
After 100 pc
dist. cut
@@
After mass
cuts
@@
Kilic et al. (2025)
(Spectroscopic SDSS sample)
@@
DESI Collaboration et al. (2025)
(Spectroscopic DESI sample)
@@ 2.1 All-Sky 100 pc Samples
We consider three Gaia defined all-sky 100 pc white dwarf samples from Early Data Release 3 and Data Release 3. The first and second, Gentile Fusillo et al. (2021) and Gaia Collaboration et al. (2021), are selected based on Gaia EDR3 photometry and astrometry, which remained unchanged in DR3. Finally, the sample of Vincent et al. (2024) uses Gaia DR3 and its new XP spectra to define an independent sample. While this sample is drawn from the full Gaia DR3, rather than existing Gaia EDR3 selections, its size and completeness are limited by the availability of Gaia XP spectroscopy. For all samples we use Gaia G, $G_{\rm BP}$, $G_{\rm RP}$, and parallax directly from Gaia DR3.
@@
For the samples of Gentile Fusillo et al. (2021) and Gaia Collaboration et al. (2021), we adopt photometric masses and $T_{\rm eff}$ from Gentile Fusillo et al. (2021), assuming hydrogen atmospheres due to the lack of consistent composition data (spectroscopy) in these samples. For the Vincent et al. (2024) sample, we adopt their photometric parameters based on Gaia XP spectra and spectral classification, which enables the use of tailored model atmospheres for the corresponding spectral classes.
Due to inaccuracies in opacities in the red wing of Lyman $\alpha$ absorption (Caron et al., 2023; O’Brien et al., 2024), the median photometric mass of white dwarfs as a function of temperature, which should remain at a constant $\approx 0.6$ M, begins to drop below $\approx 6000$ K in all samples here. To correct for this, we apply a fifth-order polynomial correction as defined by O’Brien et al. (2024) to bring the median masses back in line with $\approx 0.6$ M for all samples discussed in this work, and before any sample cuts on masses. Since photometric temperature derivations are based on the assumed mass through the mass-radius relation and the Stefan-Boltzmann law, corrections to effective temperature are also made. Since this correction uses the median masses of each sample, the exact mass and effective temperature correction applied to a star might not be the same depending on which sample it appears in. These differences are slight but in some cases cause stars to be removed from one sample but kept in another. Table 1 shows this with the column titled ‘Overlap with GF21’; all other samples are subsets of GF21, but some stars have been removed from them due to these small differences in the mass corrections.
The three all-sky samples are discussed in turn below. We also show their luminosity functions and absolute $G$ magnitude distributions in Fig. 2, compared to the 40 pc sample of O’Brien et al. (2024); Roberts et al. (2025). The absolute $G$ magnitude distributions rely solely on sample membership and Gaia DR3 data, while the luminosity functions rely on white dwarf parameters described in the last column of Table 1.
@@
With selection criteria in absolute magnitude, colour and Gaia quality flags, Gentile Fusillo et al. (2021) presents a catalogue of objects from Gaia EDR3. They calculate the probability of each object being a white dwarf based on SDSS spectroscopically confirmed white dwarfs and contaminants to provide a catalogue of $\simeq$ 359 000 high confidence white dwarf candidates ($P_{\rm WD}\geq 0.75$). Using these candidates, white dwarf masses, $T_{\rm eff}$ and $\log(g)$ were calculated using dereddened Gaia photometry for pure hydrogen, pure helium and mixed hydrogen-helium atmosphere models. Because few candidates have spectroscopic confirmation, parameters based on all three model assumptions are provided in the catalogue, along with mapped $A_{V}$ extinction.
Using the probability of being a white dwarf defined by Gentile Fusillo et al. (2021) and keeping only those objects with $P_{\rm WD}\geq 0.75$, alongside a distance cut-off at 100 pc, and limiting objects to those with a white dwarf mass between 0.541.340.54-1.34 M to eliminate double degenerate candidates and physically impossible white dwarfs if born through single star evolution, the Gentile Fusillo et al. catalogue contains 15 001 white dwarf candidates with corresponding physical parameters, which we use as a reference in Table 1 when calculating the overlap with other samples.
We note that unresolved white dwarf–main sequence binaries are already excluded to some extent in the initial Hertzsprung-Russell diagram selection of Gentile Fusillo et al. (2021). Our additional double-degenerate cut is intended to define a sample primarily composed of present-day single white dwarfs, while possible past stellar mergers are accounted for in the population synthesis (Roberts et al., 2025). Although removing unresolved binaries reduces the overall completeness of the sample, the resulting biases should remain small to moderate, since binary white dwarfs are expected to share a similar age distribution with single white dwarfs, the main parameter constrained in this work. Moreover, our analysis focuses on relative rather than absolute star formation rates.
Figure 2: Top: Luminosity function for each of the three all-sky 100 pc samples of Gentile Fusillo et al. (2021); Gaia Collaboration et al. (2021); Vincent et al. (2024) and the 40 pc sample (O’Brien et al., 2024; Roberts et al., 2025). We also include the Poisson uncertainties for each bin. Bottom: Absolute Gaia G magnitude distribution for each of the all-sky 100 pc samples and the 40 pc case, with Poisson uncertainties for each bin. The bump at $\approx$12.5 mag in the 40 pc sample does not appear in any of the 100 pc samples at the $3\sigma$ level and so appears to be an artifact of the smaller sample size rather than a real feature.

### v2

@@ Abstract
Thanks to Gaia, as well as present and future data from large-scale spectroscopic follow-up surveys (4MOST, DESI, WEAVE, SDSS-V), it is now possible to build representative and minimally biased samples of the local white dwarf population. Here we analyse several volume-limited 100 pc samples of white dwarfs, constructed from different surveys and studies, to evaluate their completeness and residual biases. We model the underlying star formation history and Galactic disc age via comparison with simulated populations of white dwarfs to quantitatively characterise completeness. We assess whether the benefit of Gaia XP spectra in datasets outweighs the reduction in sample size, and to what extent targeted, part-sky, and magnitude limited surveys can be used in comparison to all-sky volume limited surveys. Additionally, we simulate the 4MOST 100PC sub-survey and discuss its use to better understand the local star formation history.
@@
pubyear: 2025pagerange: Assessing the star formation history of all-sky and part-sky 100 pc white dwarf samplesA
@@ 1 Introduction
White dwarfs have long been recognised for their potential as chronometers, since they cool over time and have no source of nuclear fuel to maintain their temperatures (Mestel). With a close relationship between effective temperature, luminosity (or mass), and their total age (main-sequence lifetime plus white dwarf cooling time), they provide a way of rewinding time in a stellar population to look at the underlying star formation rates (Winget_1987; Oswalt_1996; Rowell_2013; Tremblay_2014; Isern_2019; Fantin_2019; Torres_2021; Cukanovaite_2023; Roberts_2025). Realising this potential, however, requires detailed models of white dwarf cooling times, main sequence lifetimes, the initial mass function, the initial-final mass relation, and a careful treatment of observational biases. These variables contribute uncertainty and cloud our knowledge of the star formation history (SFH) that produced the white dwarfs. A common benchmark is to compare such reconstructions with the formation histories inferred from main-sequence stars in the same populations (Mor_2019; Alzate_2021; DalTio_2021; Gallart_2024; Gallart_2025), with the broader goal of informing models of Galactic formation and evolution.
@@
Several methods have been used to derive the SFH of a white dwarf population based on different observational properties: estimating individual total ages for each star or using stellar population synthesis models to fit the luminosity function, absolute magnitude distribution, or Hertzsprung–Russell diagram. In a recent study, Roberts_2025 compared these approaches using the 40 pc Gaia-identified, volume-complete sample of 1073 spectroscopically confirmed white dwarfs (OBrien_2024). They found that, within systematic uncertainties and finite sample size, the different methods yielded broadly consistent results. Among them, the method based on the absolute Gaia G-band magnitude distribution was identified as the most promising, since it requires only a parallax and a G-band magnitude and can therefore be applied to larger and more distant samples with less complete follow-up data. Using these methods, they showed that a constant SFH provided a good overall fit to the sample, although variations by up to a factor of about two — either in the recent past (e.g. Mor_2019) or at earlier times—could not be ruled out given the uncertainties in population modelling.
@@
With each successive Gaia data release (Gaia_2016b; Gaia_2018; GaiaDR3_2023), the size and precision of stellar samples available for Galactic studies has grown dramatically. The advent of machine learning classification, improved stellar models and quality filtering, and increasingly realistic Galactic simulations have further enhanced our ability to extract astrophysical information from these data. In parallel, large efforts have been devoted to the spectroscopic characterisation of Gaia catalogues: from the low-resolution XP spectrophotometry provided on board Gaia, to ground-based multi-object spectroscopic (MOS) surveys such as DESI (DESI_2025), 4MOST (4MOST), WEAVE (WEAVE), and SDSS-V (SDSS-V), which together provide the medium-resolution optical ($\approx 3600-10\,000$ Å) spectroscopic context essential for Galactic archaeology and beyond. This enables more precise determinations of stellar parameters, chemical abundances, radial velocities, magnetic field strengths, and binary properties that are not possible with Gaia alone.
@@
There now exist a range of 100 pc Gaia-defined samples of confirmed and candidate white dwarfs (GentileFusillo_2021; GCNS_2021; JimenezEsteban_2023; Garcia-Zamora_2023; Vincent_2024), constructed with varying quality cuts, white dwarf probability thresholds, and each of them can be combined with auxiliary data, leading to differences in sky coverage. In particular, the footprint of Gaia-selected white dwarfs observed in MOS surveys currently covers only about 30% of the sky (DESI DR1; DESI_2025, SDSS; Kleinman2013, LAMOST; Lamost2012), with plans for each MOS mission to expand to roughly 50% of the sky by the end of operations. While these catalogues contain an order of magnitude more white dwarfs than the 40 pc sample, reducing Poisson errors on the SFH, and allowing for a wider probe of the Solar neighbourhood including the effect of the finite vertical scale height of the Galactic disc, they also lack certain things the 40 pc sample has such as either full volume completeness or spectroscopic follow up. Balancing these factors and assessing their impact is crucial for selecting the best white dwarf samples for population studies.
@@
In this work, based on the 100 pc Gaia stellar sample, we compare different white dwarf sample selections, either based on quality cuts or availability of auxiliary data such as XP spectrophotometry or MOS medium-resolution spectroscopy. We then rely on the stellar population code of Roberts_2025 to quantify these differences in terms of changes in inferred SFH and age of the Galactic disc.
@@
Identifying white dwarfs in the Gaia DR3 catalogue is challenging, even within the relatively nearby 100 pc volume. As shown for example by GentileFusillo_2021, contamination from non–white dwarf objects remains non-negligible, requiring probabilistic classification to separate white dwarfs from main-sequence stars, unresolved binaries and spurious sources. Nevertheless, the completeness of the Gaia 100 pc white dwarf sample is expected to be high, on the order of 93 per cent from a comparison with the SDSS footprint (GentileFusillo_2021). This is in line with an independent estimate by OBrien_2023; OBrien_2024, which suggests that the Gaia white dwarf selection of GentileFusillo_2021 achieves $\approx$97 per cent completeness in the 40 pc sample from a comparison with previously known white dwarfs not identified in Gaia, many of them members of unresolved binaries.
The sensitivity of the 100 pc selection is also demonstrated by different authors reaching different selections. The Gaia Collaboration’s Catalogue of Nearby Stars (GCNS_2021) (GCNS) finds 14 699 high-probability white dwarfs ($P_{\rm WD}\geq 0.75$) within 100 pc compared to the 15 001 found in GentileFusillo_2021 for the same threshold in $P_{\rm WD}$. Only 14 687 of these sources are in common between both catalogues. Also, the independent analysis of JimenezEsteban_2023 arrives at a different total of 12 718 white dwarfs within 100 pc. These reflect the different HR diagram selections, definition of white dwarf probability thresholds, and adopted quality cuts, underscoring the difficulty of assembling a complete volume census from Gaia alone.
Restricting the population to those white dwarfs with additional spectroscopic information reduces the available sample substantially. Limiting to objects with Gaia XP spectrophotometry, as explored by Vincent_2024, results in 13 235 white dwarfs within 100 pc. Limiting to objects with existing ground-based spectroscopy within the SDSS footprint (Kilic_2025_100pc) or DESI DR1 (DESI_2025) results in samples of 2 750 and 2 978 white dwarfs, respectively. While these cuts enhance the reliability of atmospheric parameter estimates and classifications, they inevitably reduce the sample size and may introduce biases. Most of this loss arises from the limited (primarily northern hemisphere) footprints of these MOS surveys, but within their surveyed conical 100 pc volumes the resulting samples remain reasonably complete and representative (Kilic_2020; Manser_2024).
@@
Table 1: Sample sizes and quality cuts made for the samples considered in this work. The GCNS catalogue (GCNS_2021) also includes main-sequence stars, hence the large number of objects removed after the $P_{\rm WD}$ white dwarf selection. The Jiménez-Esteban sample does not have masses and effective temperatures for all stars in the sample; where they are missing, the values from GF21 are used instead. The part-sky DESI and SDSS footprints are shown in Fig. 3. ‘tailored atm’ implies that photometric or spectroscopic fits were performed with model atmospheres using an appropriate chemical composition for a given known spectral type. In all cases, photometric atmospheric parameters were corrected for the low-mass problem at $T_{\rm eff}<6000$ K (OBrien_2024) before the mass cuts for binary removal were performed. Note that in the absolute $G$ magnitude distribution method, we do not use the derived stellar parameters.
@@
Stars within
100 pc in sample
@@
After
mass cuts
@@
Kilic_2025_100pc
(Spectro. SDSS sample)
@@
DESI_2025 + this work
(Spectro. DESI sample)
@@ 2.1 All-Sky 100 pc Samples
We consider three Gaia defined all-sky 100 pc white dwarf samples from Early Data Release 3 and Data Release 3. The first and second, GentileFusillo_2021 and GCNS_2021, are selected based on Gaia EDR3 photometry and astrometry, which remained unchanged in DR3. Finally, the sample of Vincent_2024 uses Gaia DR3 and its new XP spectra to define an independent sample. While this sample is drawn from the full Gaia DR3, rather than existing Gaia EDR3 selections, its size and completeness are limited by the availability of Gaia XP spectroscopy. For all samples we use Gaia G, $G_{\rm BP}$, $G_{\rm RP}$, and parallax directly from Gaia DR3.
@@
For the samples of GentileFusillo_2021 and GCNS_2021, we adopt photometric masses and $T_{\rm eff}$ from GentileFusillo_2021, assuming hydrogen atmospheres due to the lack of consistent composition data (spectroscopy) in these samples. For the Vincent_2024 sample, we adopt their photometric parameters based on Gaia XP spectra and spectral classification, which enables the use of tailored model atmospheres for the corresponding spectral classes.
Due to inaccuracies in opacities in the red wing of Lyman $\alpha$ absorption (Caron_2023; OBrien_2024; Sahu_2025), the median photometric mass of white dwarfs as a function of temperature, which should remain at a constant $\approx 0.6$ M, begins to drop below $\approx 6000$ K in all samples here. To correct for this, we apply a fifth-order polynomial correction as defined by OBrien_2024 to bring the median masses back in line with $\approx 0.6$ M for all samples discussed in this work, and before any sample cuts on masses. Since photometric temperature derivations are based on the assumed mass through the mass-radius relation and the Stefan-Boltzmann law, corrections to effective temperature are also made. Since this correction uses the median masses of each sample, the exact mass and effective temperature correction applied to a star might not be the same depending on which sample it appears in. These differences are slight but in some cases cause stars to be removed from one sample but kept in another. Table 1 shows this with the column titled ‘Overlap with GF21’; all other samples are subsets of GF21, but some stars have been removed from them due to these small differences in the mass corrections.
The three all-sky samples are discussed in turn below. We also show their luminosity functions and absolute $G$ magnitude distributions in Fig. 2, compared to the 40 pc sample of OBrien_2024; Roberts_2025. The absolute $G$ magnitude distributions rely solely on sample membership and Gaia DR3 data, while the luminosity functions rely on white dwarf parameters described in the last column of Table 1 along with the mass-radius relation of Bedard_2020.
@@
With selection criteria in absolute magnitude, colour and Gaia quality flags, GentileFusillo_2021 presents a catalogue of objects from Gaia EDR3. They calculate the probability of each object being a white dwarf based on SDSS spectroscopically confirmed white dwarfs and contaminants to provide a catalogue of $\simeq$ 359 000 high confidence white dwarf candidates ($P_{\rm WD}\geq 0.75$). Using these candidates, white dwarf masses, $T_{\rm eff}$ and $\log(g)$ were calculated using dereddened Gaia photometry for pure hydrogen, pure helium and mixed hydrogen-helium atmosphere models. Because few candidates have spectroscopic confirmation, parameters based on all three model assumptions are provided in the catalogue, along with mapped $A_{V}$ extinction.
Using the probability of being a white dwarf defined by GentileFusillo_2021 and keeping only those objects with $P_{\rm WD}\geq 0.75$, alongside a distance cut-off at 100 pc, and limiting objects to those with a white dwarf mass between 0.541.340.54-1.34 M to eliminate double degenerate candidates and physically impossible white dwarfs if born through single star evolution, the Gentile Fusillo et al. catalogue contains 15 001 white dwarf candidates with corresponding physical parameters, which we use as a reference in Table 1 when calculating the overlap with other samples. While many factors such as metallicity, rotation, magnetic effects and more can produce lower mass white dwarfs through single star evolution (Renedo_2010; Cunningham_2020), these are not currently fully included in empirical IFMRs. As an approximation, we therefore use the hard cut off of 0.54 $M_{\odot}$. The upper limit is to enforce the Chandrasekhar mass limit for white dwarfs (Chandrasekhar_1931).
We note that unresolved white dwarf–main sequence binaries are already excluded to some extent in the initial Hertzsprung-Russell diagram selection of GentileFusillo_2021. Our additional double-degenerate cut is intended to define a sample primarily composed of present-day single white dwarfs, while possible past stellar mergers are accounted for in the population synthesis (Roberts_2025). Although removing unresolved binaries reduces the overall completeness of the sample, the resulting biases should remain small to moderate, since binary white dwarfs are expected to share a similar age distribution with single white dwarfs, the main parameter constrained in this work. Moreover, our analysis focuses on relative rather than absolute star formation rates.
Figure 2: Top: Luminosity function for each of the three all-sky 100 pc samples of GentileFusillo_2021; GCNS_2021; Vincent_2024 and the 40 pc sample (OBrien_2024; Roberts_2025). We also include the Poisson uncertainties for each bin. Bottom: Absolute Gaia G magnitude distribution for each of the all-sky 100 pc samples and the 40 pc case, with Poisson uncertainties for each bin. The bump at $\approx$12.5 mag in the 40 pc sample does not appear in any of the 100 pc samples at the $3\sigma$ level and so appears to be an artifact of the smaller sample size rather than a real feature.